OpenStudy (anonymous):

find f with an exponent of -1 for f(x)= -3x+4

OpenStudy (anonymous):

this means the inverse function yes?

OpenStudy (anonymous):

your function says multiply by -3 then add 4. your inverse will say subtract 4 then divide by -3 it is written as \[f^{-x}(x)=\frac{x-4}{-3}\]

OpenStudy (anonymous):

or you can write \[y=-3x+4\] switch x and y to get \[x=-3y+4\] and solve for x you get \[x-4=-3y\] \[\frac{x-4}{-3}=y\] so \[f^{-1}(x)=\frac{x-4}{-3}=\frac{4-x}{3}\]

OpenStudy (anonymous):

sorry, i am writing to make sure i got it

OpenStudy (anonymous):

this is exactly the same as saying \[-3x+4=13\] solve for x. what would you do? subtract 4, divide by -3. that is what the inverse function accomplishes

OpenStudy (anonymous):

wow! so 4-x/3 is the answer